Use the Substitution method to solve the system of equations.
x + y = -4
x - y = 2

Answer by checkley77(12844)

x + y = -4
x=-4-y
x - y = 2
(-4-y)-y=2
-4-y-y=2
-2y=2+4
-2y=6
y=6/-2
y=-3 answer.
x+(-3)=-4
x-3=-4
x=-4+3
x=-1 answer.
proof:
-1-(-3)=2
-1+3=2
2=2